OpenStudy (anonymous):

Solve the differential equation(antiderivatives calculus 2) f(x)=  12x^3+ 6x^2-7 with initial condition f(1)=  4

OpenStudy (amistre64):

Hi Saif, you here?

OpenStudy (anonymous):

yea i am here

OpenStudy (anonymous):

its college calculus 2, antiderivatives are killin me. SO confusing

OpenStudy (anonymous):

Confused. There is no differential as written. Does "dx" multiply the polynomial?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

Are you asking us to just integrate/ find the antiderivative of the function f(x)?

OpenStudy (amistre64):

You want to the integral and its anchor at (1,4) right?

OpenStudy (anonymous):

yea i think so I thought it was just asking the antiderivative with that condition

OpenStudy (amistre64):

F(x)= (12/4)x^4+ (6/3)x^3 -7x + C F(x) = 3x^4 +2x^3 -7x +C solve for F(1) = 4 :)

OpenStudy (anonymous):

you are using big F(x), its giving you f'(x), wouldnt the notation change?

OpenStudy (anonymous):

its not the same convention as a general antiderivative

OpenStudy (amistre64):

naming is naming, I just did it to be flamboyant :)

OpenStudy (anonymous):

ok so the totation would be f(x) correct and then you find f(1)

OpenStudy (anonymous):

i mean you find C

OpenStudy (amistre64):

We find the "antiderivative" of the function to find a family of curves that we can adopt... then we apply (1,4) into it to see what our anchoring constant turns out to be, right?

OpenStudy (anonymous):

yea correct

OpenStudy (amistre64):

then "suit up" and find our constant :)

OpenStudy (anonymous):

i got 6

OpenStudy (amistre64):

4 = 3 +2 -7 +C 4 = 5-7 + C 4 = -2 + C 6 = C F(x) = 3x^4 +2x^3 -7x +6 yay!!!

OpenStudy (anonymous):

thanks so much. Can you take a look at the other antiderivative I posted. I am getting confused, he took like 2 antiderivatives?

OpenStudy (amistre64):

Didja notice me new "title"..... funny, i never was a sandwich before :)

OpenStudy (anonymous):

hahaahahaah

OpenStudy (amistre64):

ill check em out...

OpenStudy (anonymous):

its 2 below the this question